\begin{thebibliography}{99}
\bibitem{Bakh} Bakhvalov, A. N. Continuity in $\Lambda $-variation of
functions of several variables and the convergence of multiple Fourier
series (Russian). \textit{Mat. Sb.} 193, 12(2002), 3--20; English transl.in
\textit{Sb. Math.} 193(2002), No. 11-12, 1731--1748.

\bibitem{Bakh2} Bakhvalov, A. N. Continuity in $\Lambda $-variation and the
summation of multiple Fourier series by Ces\`{a}ro methods. (Russian) Mat.
Zametki 90(2011), No. 4, 483--500; translation in Math. Notes 90(2011),
No. 3-4, 469--484.

\bibitem{Bakh3} Bakhvalov, A. N. On the localization of Ces\`{a}ro means of
Fourier series of functions of bounded $\Lambda $-variation. (Russian) Izv.
Vyssh. Uchebn. Zaved. Mat. 55(2011), , no. 8, 9--13; translation in Russian
Math. (Iz. VUZ) 55 (2011), No. 8, 7--10.

\bibitem{Ch} Chanturia, Z. A, The modulus of variation of a function and its
application in the theory of Fourier series, \textit{Soviet. Math. Dokl.} 15(1974), 67-71.

\bibitem{D1} Dyachenko M. I, Waterman classes and spherical partial sums of
double Fourier series, Anal. Math. 21(1995), 3-21.

\bibitem{D2} Dyachenko M. I, Two-dimensional Waterman classes and $u$%
-convergence of Fourier series (Russian).\textit{Mat. Sb.} 190(1999), No. 7,
23--40; English transl.in \textit{Sb. Math.} 190(1999), No. 7-8, 955--972.

\bibitem{DW} Dyachenko M. I.; Waterman D. , Convergence of double Fourier
series and W-classes, Trans. Amer. Math. Soc. 357(2005), 397-407.

\bibitem{GoPRI} Goginava U. On the uniform summability of two-dimensional
trigonometric Fourier series. Proc. A. Razmadze Math. Inst. 124 (2000),
55--72.

\bibitem{GoJAT} Goginava U., Uniform convergence of Ces\'{a}ro means of
negative order of double Walsh-Fourier series. \textit{J. Approx. Theory. }
 124(2003), 96-108.

\bibitem{GoEJA} Goginava U., On the uniform convergence of multiple
trigonometric Fourier series. \textit{East J. Approx.} 5(1999), No. 3, 253-266.

\bibitem{EJA} Goginava U., Sahakian A, On the convergence of Fourier series
of functions of bounded partial generalized variation, East J. Approx. 16(2010), No. 2, 153--165.


\bibitem{SMJ} Goginava U., Cesaro means of negative order of double Fourier
series and Generalized Bounded variation, Siberian Mathematical Journal,
54(2013), , No. 6, pp. 1005--1013.

\bibitem{Szeg} Goginava U., Sahakian A, On the convergence of Ces\'aro means
of negative order of double trigonometric Fourier series of functions of
bounded partial generalized variation, Acta. Sci. Math. (Szeged) 77(2011),
451-471.

\bibitem{GMJ} Goginava U., Sahakian A, Convergence of double Fourier series
and generalized $\Lambda$-variation, Georgian Math. J. 19(2012), No. 3, 497-509.

%\bibitem{CMA} Goginava U., Sahakian A, On the convergence of multiple Walsh-Fourier series of functions of bounded generalized variation, Journal of Contemporary Mathematical Analysis 5, 47 (2012), 221-233.

\bibitem{AM} Goginava U., Sahakian A, On the convergence of multiple Fourier
series of functions of bounded partial generalized variation, Anal. Math. 39(2013), No. 1, 45--56.

\bibitem{Stek} Goginava U., Sahakian A, On the summability of
multipleFourier series of functions of bounded partial generalized variation
Proc. Steklov Inst. Math, 280(2013), 150-161.


\bibitem{CMA1} Goginava U., Sahakian A, On the Convergence and Summability
of double Walsh-Fourier series of functions of bounded generalized variation.
Journal of Contemporary Mathematical Analysis  48(2014), No. 6.

\bibitem{SBM} Goginava U., Sahakian A, Convergence of Multiple Fourier Series of Functions of Bounded Generalized Variation (Submitted)

\bibitem{Ha} Hardy G. H., On double Fourier series and especially which
represent the double zeta function with real and incommensurable parameters.
\textit{Quart. J. Math. Oxford Ser.} {37}(1906), 53-79.

\bibitem{Jo} Jordan C., Sur la series de Fourier. \textit{C.R. Acad. Sci.
Paris}. {92}(1881), 228-230.

\bibitem{M} Marcinkiewicz J. On a class of functions and their Fourier
series. \textit{Compt. Rend. Soc. Sci. Warsowie}, 26(1934), 71-77.

\bibitem{Sab} Sablin A. I., $\Lambda $-variation and Fourier series
(Russian), \textit{Izv. Vyssh. Uchebn. Zaved. Mat.} 10(1987), 66--68;
English transl. in \textit{\ Soviet Math. (Iz. VUZ)} 31(1987).

\bibitem{SG} Sargsyan O. G., On the convergence and the Gibbs phenomenon of
multiple Fourier series for functions of bounded harmonic variation
(Russian). \textit{Izv. Akad. Nauk Armenii Mat.}, 28(1993), No. 3, 3-20, English
transl. in \textit{J. Contemp. Math. Anal.}, 28(1993), No. 3.

\bibitem{Saha} Sahakian A. A., On the convergence of double Fourier series
of functions of bounded harmonic variation (Russian). \textit{\ Izv. Akad.
Nauk Armyan. SSR Ser. Mat.} 21(1986), No. 6, 517-529; English transl., \textit{\
Soviet J. Contemp. Math. Anal.} 21(1986), No. 6, 1-13.

\bibitem{Wa2} Waterman D., On the summability of Fourier series of functions
of $\Lambda $-bounded variation, Studia Math. 55(1976), 87-95.


\bibitem{W} Waterman D., On convergence of Fourier series of functions of
generalized bounded variation. \textit{Studia Math.}, 44(1972), No. 1, 107-117.

\bibitem{Wi} Wiener N., The quadratic variation of a function and its
Fourier coefficients. \textit{Massachusetts J. Math}., 3(1924), 72-94.

\bibitem{Zh} Zhizhiashvili L. V., Trigonometric Fourier series and their
conjugates, Kluwer Academic Publishers, Dobrecht, Boston, London, 1996.

\bibitem{Zy} Zygmund A., Trigonometric series. Cambridge University Press,
Cambridge, 1959.
\end{thebibliography}

\documentclass[11pt]{amsart}
\usepackage{amsmath}
\usepackage{amsfonts}
\usepackage{amsthm}
\usepackage[active]{srcltx}

\newtheorem{theorem}{Theorem}[section]

\newtheorem{remark}{Remark}[section]
\newtheorem{definition}{Definition}[section]
\newtheorem{corollary}{Corollary}[section]
\newtheorem*{TDW}{Theorem DW}

\numberwithin{equation}{section}

\newtheorem*{Go2}{Theorem G2}

\newtheorem*{Wa2}{Theorem W2}
\newtheorem*{Sa}{Theorem S}
\newtheorem*{Zh}{Theorem Zh}

\newtheorem*{Bakh1}{Theorem B1}
\newtheorem*{Bakh2}{Theorem B2}

\begin{document}
\author{Ushangi Goginava and Artur Sahakian}
\title[convergence and Summability of Fourier series ]{Convergence and
Summability of Multiple Fourier series and generalized variation}
\address{U. Goginava, Institute of Mathematics, Faculty of Exact and Natural
Sciences, Tbilisi State University, Chavchavadze str. 1, Tbilisi 0128,
Georgia}
\email{zazagoginava@gmail.com}
\address{A. Sahakian, Yerevan State University, Faculty of Mathematics and
Mechanics, Alex Manoukian str. 1, Yerevan 0025, Armenia}
\email{sart@ysu.am}
\maketitle

\begin{abstract}
In this paper we present results on convergence and Ces\`{a}ro summability of Multiple Fourier series of functions of bounded generalized variation.
\end{abstract}

\footnotetext{%
2010 Mathematics Subject classification: 26A45
\par
Key words and phrases: Waterman's class, generalized Wiener's class,
Multiple Fourier series, Cesaro means
\par
The reseach of U. Goginava was supported by Shota Rustaveli National Science
Foundation grant no.13/06 (Geometry of function spaces, interpolation and
embedding theorems).}

\section{Classes of Functions of two variables of Bounded Generalized
Variation}

In 1881 Jordan \cite{Jo} introduced a class of functions of bounded
variation and applied it to the theory of Fourier series. This
notion was generalized hereinafter by many authors (quadratic variation, $\Phi$%
-variation, $\Lambda$-variation ets., see \cite{Wi}-\cite{Ch}). In two
dimensional case the class BV of functions of bounded variation was
introduced by Hardy \cite{Ha}.

In this section we introduce several classes of bivariate functions of Bounded Generalized Variation  and compare them with the class $HBV$ (se Definition \ref{def1.1} below), which is important for the applications in Fourier analysis (see Theorem S in Section 2.).

Let $f(x,y),\ (x,y)\in\mathbb{R}^2$ be a real function of two variables of period $2\pi $ with respect to
each variable. Given intervals $I=(a,b)$, $J=(c,d)$ and points $x,y$ from $%
T:=[0,2\pi ]$ we denote
\begin{equation*}
f(I,y):=f(b,y)-f(a,y),\qquad f(x,J)=f(x,d)-f(x,c)
\end{equation*}
and
\begin{equation*}
f(I,J):=f(a,c)-f(a,d)-f(b,c)+f(b,d).
\end{equation*}
Let $E=\{I_{i}\}$ be a collection of nonoverlapping intervals from $T$
ordered in arbitrary way and let $\Omega $ be the set of all such
collections $E$. Denote by $\Omega _{n}$ set of all collections of $n$
nonoverlapping intervals $I_{k}\subset T.$

For the sequence of positive numbers $\Lambda =\{\lambda
_{n}\}_{n=1}^{\infty }$ we define
\begin{equation*}
\Lambda V_{1}(f)=\sup_{y}\sup_{E\in \Omega }\sum_{n}\frac{|f(I_{i},y)|}{%
\lambda _{i}}\,\,\,\,\,\,\left( E=\{I_{i}\}\right) ,
\end{equation*}
\begin{equation*}
\Lambda V_{2}(f)=\sup_{x}\sup_{F\in \Omega }\sum_{m}\frac{|f(x,J_{j})|}{%
\lambda _{j}}\qquad (F=\{J_{j}\}),
\end{equation*}
\begin{equation*}
\Lambda V_{1,2}(f)=\sup_{F,\,E\in \Omega }\sum_{i}\sum_{j}\frac{%
|f(I_{i},J_{j})|}{\lambda _{i}\lambda _{j}}.
\end{equation*}

\begin{definition}\label {def1.1}
We say that the function $f$ has Bounded $\Lambda $-variation on $%
T^{2}=[0,2\pi ]^{2}$ and write $f\in \Lambda BV$, if
\begin{equation*}
\Lambda V(f):=\Lambda V_{1}(f)+\Lambda V_{2}(f)+\Lambda V_{1,2}(f)<\infty .
\end{equation*}%
We say that  $f$ has Bounded Partial $\Lambda $-variation and
write $f\in P\Lambda BV$ if
\begin{equation*}
P\Lambda V(f):=\Lambda V_{1}(f)+\Lambda V_{2}(f)<\infty .
\end{equation*}
\end{definition}

If $\lambda_n\equiv 1$ (or if $0<c<\lambda_n<C<\infty,\ n=1,2,\ldots$) the
classes $\Lambda BV$ and $P\Lambda BV$ coincide with the Hardy class $BV$
and PBV respectively. Hence it is reasonable to assume that $%
\lambda_n\to\infty$ and since the intervals in $E=\{I_i\}$ are ordered
arbitrarily, we will suppose, without loss of generality, that the sequence $%
\{\lambda_n\}$ is increasing. Thus,
\begin{equation}  \label{Lambda}
1<\lambda_1\leq \lambda_2\leq\ldots,\qquad \lim_{n\to\infty}\lambda_n=\infty.
\end{equation}

In the case when $\lambda _{n}=n,\ n=1,2\ldots $ we say \textit{Harmonic
Variation} instead of $\Lambda $-variation and write $H$ instead of $\Lambda$
($HBV$, $PHBV$, $HV(f)$, ets).

The notion of $\Lambda$-variation was introduced by D. Waterman \cite{W} in
one dimensional case and A. Sahakian \cite{Saha} in two dimensional case. The class $PBV$ as well as the class $PBV_{p}$ (see Definition \ref{def1.2}) was introdused by U. Goginava in \cite{GoEJA}.

\begin{definition}\label{def1.2}
Let $\Phi $-be a strictly increasing continuous function on $[0,+\infty )$
with $\Phi \left( 0\right) =0$. We say that the function $f$ has Bounded
Partial $\Phi $-variation on $T^{2}$ and write $f\in PBV_\Phi$, if
\begin{equation*}
V_{\Phi }^{\left( 1\right) }\left( f\right)
:=\sup\limits_{y}\sup\limits_{\{I_{i}\}\in \Omega
_{n}}\sum\limits_{i=1}^{n}\Phi \left( |f\left( I_{i},y\right) |\right)
<\infty ,\quad n=1,2,...,
\end{equation*}
\begin{equation*}
V_{\Phi }^{\left( 2\right) }\left( f\right)
:=\sup\limits_{x}\sup\limits_{\{J_{j}\}\in \Omega
_{m}}\sum\limits_{j=1}^{m}\Phi \left( |f\left( x,J_{j}\right) |\right)
<\infty ,\quad m=1,2,....
\end{equation*}
In the case when $\Phi \left( u\right) =u^{p},\,p\geq 1$, we say that $f$ has Bounded Partial $p$-variation and write $f\in PBV_{p}$.
\end{definition}
In the following theorem the necessary and sufficient conditions are obtained for the inclusion $P\Lambda BV\subset HBV$.
\begin{theorem}[U. Goginava, A. Sahakian \cite{EJA}]\label{T1}
Let $\Lambda =\{\lambda _{n}\}$ with $\lambda _{n}=n\gamma_n$ and $\gamma_{n}\geq
\gamma _{n+1}>0,\ n=1,2,....\,\,\,$. \newline
\indent
1) If
\begin{equation}  \label{T1-1}
\sum_{n=1}^{\infty }\frac{\gamma _{n}}{n}<\infty,
\end{equation}
\hskip10mm then $P\Lambda BV\subset HBV$.\newline
\indent 2) If  $\gamma_n=O(\gamma_{n^{[1+\delta]}})$ for some $\delta>0$ and
\begin{equation*}
\sum_{n=1}^{\infty }\frac{\gamma _{n}}{n}=\infty,
\end{equation*}
\hskip10mm then $P\Lambda BV\not\subset HBV$.
\end{theorem}

\begin{corollary}\label{cor1.1}
$PBV\subset HBV$ and $PHBV\not\subset HBV$.
\end{corollary}
\begin{corollary}\label{cor1.2}
Let $\Phi $ and $\Psi $ are conjugate functions in the sense of Yung ($%
ab\leq \Phi (a)+\Psi (b)$) and let for some $\{\lambda _{n}\}$ satisfying  (\ref{Lambda}),
\begin{equation}
\sum_{n=1}^{\infty }\Psi \left( \frac{1}{\lambda _{n}}\right) <\infty .
\label{Phi}
\end{equation}%
Then $PBV_{\Phi }\subset HBV$. In particular, $PBV_{p}\subset HBV$ for any $%
p>1$.
\end{corollary}

\begin{definition}[U. Goginava  \protect\cite{GoEJA}]
The Partial Modulus of Variation of a function $f$ are the functions $%
v_{1}\left( n,f\right)$ and $v_{2}\left( m,f\right) $ defined by
\begin{eqnarray*}
&&
v_{1}\left( n,f\right) :=\sup\limits_{y}\sup\limits_{\{I_{i}\}\in \Omega
_{n}}\sum\limits_{i=1}^{n}\left| f\left( I_{i},y\right) \right| ,\quad
n=1,2,\ldots,
\\&&
v_{2}\left( m,f\right) :=\sup\limits_{x}\sup\limits_{\{J_{k}\}\in \Omega
_{m}}\sum\limits_{i=1}^{m}\left| f\left( x,J_{k}\right) \right| ,\quad
m=1,2,\ldots.
\end{eqnarray*}
\end{definition}
For functions of one variable the concept of the modulus of variation was
introduced by Chanturia \cite{Ch}.
\begin{theorem}[U. Goginava, A. Sahakian \cite{EJA}]\label{T2}
Let $f$ be such that
\begin{equation*}
\sum\limits_{n=1}^{\infty }\frac{\sqrt{v_{j}\left( n,f\right) }}{n^{3/2}}%
<\infty ,\quad j=1,2.
\end{equation*}%
Then $f\in HBV.$
\end{theorem}
Another class of functions of generalized bounded variation was introduced by
M. Dyachenko and D. Waterman in \cite{DW}. Denoting by $\Gamma $ the the set of finite
collections of nonoverlapping rectangles $A_{k}:=\left[ \alpha _{k},\beta
_{k}\right] \times \left[ \gamma _{k},\delta _{k}\right] \subset T^{2}$ they
define
\begin{equation*}
\Lambda ^{\ast }V\left( f\right) :=\sup_{\{A_{k}\}\in \Gamma }\sum\limits_{k}%
\frac{\left\vert f\left( A_{k}\right) \right\vert }{\lambda _{k}}.
\end{equation*}

\begin{definition}[M. Dyachenko, D. Waterman \cite{DW}]
 We say that $f\in \Lambda ^{\ast }BV $ if
\begin{equation*}
\Lambda V(f):=\Lambda V_{1}(f)+\Lambda V_{2}(f)+\Lambda ^{\ast }V\left(
f\right)<\infty .
\end{equation*}
\end{definition}

In \cite{GMJ} we introduced a new classes of functions of generalized
bounded variation and investigate the convergence of Fourier series of
function of that classes.
For the sequence $\Lambda =\{\lambda _{n}\}_{n=1}^{\infty }$ we define
\begin{equation*}
\Lambda ^{\#}V_{1}(f)=\sup_{\{y_{i}\}\subset T}\sup_{\{I_{i}\}\in \Omega
}\sum_{i}\frac{|f(I_{i},{y_{i}})|}{\lambda _{i}},
\end{equation*}%
\begin{equation*}
\Lambda ^{\#}V_{2}(f)=\sup_{\{x_{j}\}\subset T}\sup_{\{J_{j}\}\in \Omega
}\sum_{j}\frac{|f(x_{j},J_{j}|}{\lambda _{j}}.
\end{equation*}
\begin{definition}[U. Goginava, A. Sahakian \cite{EJA}]
We say that $f\in\Lambda ^{\#}BV$, if
\begin{equation*}
\Lambda ^{\#}V(f):=\Lambda ^{\#}V_{1}(f)+\Lambda ^{\#}V_{2}(f)<\infty .
\end{equation*}
\end{definition}
It is not hard to see, that
\begin{equation}
\Lambda ^{\ast }BV\subset \Lambda ^{\#}BV\subset P\Lambda BV.  \label{cond0}
\end{equation}%
Obviously,  the function $f(x,y)=\mathrm{sign}(x-y)$ belongs
to  $P\Lambda BV\setminus \Lambda ^{\#}BV$ for any $\Lambda $. On the
other hand, we have proved the following result.
\begin{theorem}[U. Goginava, A. Sahakian \cite{GMJ}]
If $\Lambda =\{\lambda _{n}\}$ and%
\begin{equation*}
{\limsup\limits_{n\rightarrow \infty }}\left( \sum\limits_{k=1}^{n^{2}}%
\frac{1}{\lambda _{k}}\right)  \left( \sum\limits_{k=1}^{n}\frac{1}{%
\lambda _{k}}\right)^{-1} =+\infty ,
\end{equation*}
then $\Lambda ^{\#}BV\setminus \Lambda ^{\ast }BV\neq \emptyset $.
\end{theorem}
In the next theorem we characterize sequences  $\Lambda =\{\lambda _{n}\}$ for which the inclusion
$\Lambda ^{\#}BV\subset HBV$ holds.
\begin{theorem}[U. Goginava, A. Sahakian \cite{GMJ}]\label{TT}
Let $\Lambda =\{\lambda _{n}\}$.\newline
\indent a) If
\begin{equation*}
{\limsup\limits_{n\rightarrow \infty }}\frac{\lambda _{n}\log n}{n}<\infty ,
\end{equation*}%
\hskip10mm then%
\begin{equation*}
\Lambda ^{\#}BV\subset HBV.
\end{equation*}%
\indent b) If $\frac{\lambda _{n}}{n}\downarrow 0$ and%
\begin{equation*}
{\limsup\limits_{n\rightarrow \infty }}\frac{\lambda _{n}\log n }{n}=+\infty ,
\end{equation*}%
\hskip10mm  then%
\begin{equation*}
\Lambda ^{\#}BV\not\subset HBV.
\end{equation*}%
\end{theorem}

\begin{definition}[U. Goginava, A. Sahakian \cite{GMJ}]
Let $\Phi $-be a strictly increasing continuous function on $[0,+\infty )$
with $\Phi \left( 0\right) =0$. We say that the function $f\in B^{\#}V_{\Phi
}\left( T^{2}\right) $, if
\begin{equation*}
V_{\Phi ,1}^{\#}\left( f\right) :=\sup\limits_{\{y_{i}\}\subset
T}\sup\limits_{\{I_{i}\}\in \Omega }\sum\limits_{i}\Phi \left( |f\left(
I_{i},y_{i}\right) |\right) <\infty ,
\end{equation*}%
and%
\begin{equation*}
V_{\Phi ,2}^{\#}\left( f\right) :=\sup\limits_{\{x_{j}\}\subset
T}\sup\limits_{\{J_{j}\}\in \Omega }\sum\limits_{j}\Phi \left( |f\left(
x_{j},J_{j}\right) |\right) <\infty .
\end{equation*}
\end{definition}

Next, we define
\begin{equation*}
v_{1}^{\#}\left( n,f\right)
:=\sup\limits_{\{y_{i}\}_{i=1}^{n}}\sup\limits_{\{I_{i}\}\in \Omega
_{n}}\sum\limits_{i=1}^{n}\left\vert f\left( I_{i},y_{i}\right) \right\vert
,\quad n=1,2,\ldots ,
\end{equation*}%
\begin{equation*}
v_{2}^{\#}\left( m,f\right)
:=\sup\limits_{\{x_{j}\}_{j=1}^{m}}\sup\limits_{\{J_{k}\}\in \Omega
_{m}}\sum\limits_{j=1}^{m}\left\vert f\left( x_{j},J_{j}\right) \right\vert
,\quad m=1,2,\ldots .
\end{equation*}

\begin{theorem}[U. Goginava, A. Sahakian \cite{GMJ}]\label{phi}
Let $\Phi $ and $\Psi $ are conjugate functions in the sense of
Yung ($ab\leq \Phi (a)+\Psi (b)$) and let
\begin{equation*}
\sum_{n=1}^{\infty }\Psi\left(  \frac{\log n}n\right)
<\infty .
\end{equation*}%
Then%
\begin{equation*}
B^{\#}V_{\Phi }\subset \left\{ \frac{n}{\log n }\right\}
^{\#}BV.
\end{equation*}
\end{theorem}

\begin{theorem}[U. Goginava, A. Sahakian \cite{GMJ}]
\label{v(n)}Let%
\begin{equation*}
\sum\limits_{n=1}^{\infty }\frac{v_{s}^{\#}\left( f,n\right) \log n }{n^{2}}<\infty ,\qquad \ s=1,2.
\end{equation*}%
Then
$$f\in \left\{ \frac{n}{\log n }\right\} ^{\#}BV.$$
\end{theorem}
Observe that by Theorem \ref{TT} we have the inclusion $\left\{ \frac{n}{\log n }\right\} ^{\#}BV\subset HBV$.
Now, for a sequence $\Lambda=\{\lambda_n\}$ we denote
$$
\Lambda _{n}:=\left\{ \lambda _{k}\right\} _{k=n}^{\infty },\qquad n=1,2,\ldots
$$


\begin{definition}[U Goginava \cite{SMJ}]
We say that the function $f\in \Lambda^{\#} BV$\ is continuous in $\Lambda ^{\#}$-variation
and write $f\in C\Lambda ^{\#}V$, if%
\begin{equation*}
\lim\limits_{n\rightarrow \infty }\Lambda _{n}^{\#}V_{1}\left( f\right)
=\lim\limits_{n\rightarrow \infty }\Lambda _{n}^{\#}V_{2}\left( f\right) =0.
\end{equation*}%
\end{definition}

\begin{theorem}[U. Goginava, A. Sahakian \cite {CMA1}]
\label{continous} Let the sequence $\Lambda =\left\{ \lambda _{n}\right\} $
be such that
\begin{equation*}\label{maincond}
\liminf_{n\rightarrow \infty }\frac{\lambda _{2n}}{\lambda _{n}}=q>1.
\end{equation*}%
Then $\Lambda ^{\#}BV=C\Lambda ^{\#}V$.
\end{theorem}

\begin{theorem}[U. Goginava \cite{SMJ}]
\label{MV}Let $\alpha +\beta <1,\alpha ,\beta >0$ and%
\begin{equation*}
\sum\limits_{j=1}^{\infty }\frac{v_{s}^{\#}\left( f;2^{j}\right) }{%
2^{j\left( 1-\left( \alpha +\beta \right) \right) }}<\infty ,\qquad s=1,2.
\end{equation*}%
Then $f\in C\{n^{1-\left( \alpha +\beta \right) }\}^{\#}V$.
\end{theorem}

\section{Convergence of double Fourier series}
Everywhere in this and in the next section we suppose that the function $f$ is measurable on $\mathbb R^2$ and $2\pi $-periodic  with respect to each variable. The double Fourier series of a function  $f\in L^{1}\left( T^{2}\right)$ with respect to the trigonometric system is the series
\begin{equation*}
S \left[ f\right] :=\sum_{m,n=-\infty }^{+\infty }\widehat{f}\left(
m,n\right) e^{imx}e^{iny},
\end{equation*}
where
\begin{equation*}
\widehat{f}\left( m,n\right) =\frac{1}{4\pi ^{2}}\int_{0}^{2\pi
}\int_{0}^{2\pi }f(x,y)e^{-imx}e^{-iny}dxdy
\end{equation*}
are the Fourier coefficients of  $f$. The rectangular partial
sums of $S[f]$ are defined as follows:
\begin{equation*}
S_{M,N} \left[ f, (x,y)\right] :=\sum_{m=-M } ^{M } \sum_{n=-N }^N \widehat{f%
}\left( m,n\right) e^{imx}e^{iny},
\end{equation*}
In this paper we consider only {\bf  Pringsheim convergence } of double Fourier series, i.e {\bf convergence of rectangular partial
sums}   $S_{M,N} \left[ f, (x,y)\right]$, as $M,N\to\infty$.

We denote by $C(T^{2})$ the space of continuous on $\mathbb R^2$  and $2\pi $%
-periodic with respect to each variable  functions with the norm
\begin{equation*}
\Vert f\Vert _{C}:=\sup_{x,y\in T^{2}}|f(x,y)|.
\end{equation*}

For a function $f$ we denote by $f\left( x\pm 0,y\pm
0\right) $ the open coordinate quadrant limits (if exist) at the point $%
\left( x,y\right) $ and let $f^{\ast }(x,y)$
be the arithmetic mean of that quadrant limits:
\begin{multline}\label{limits}
f^{\ast }(x,y) :=\frac{1}{4}\left\{ f\left( x+0,y+0\right) +f\left(
x+0,y-0\right) \right.  \\
\left. +f\left( x-0,y+0\right) +f\left( x-0,y-0\right) \right\} .
\end{multline}
\begin{remark}
Observe that for a function $f\in \Lambda BV$ the quadrant limits $f\left( x\pm 0,y\pm 0\right) $
may not exist. As was shown in \cite{GMJ} for any function $f\in \Lambda ^{\#}BV$ the quadrant limits $f\left( x\pm 0,y\pm 0\right) $ exist at any point $\left( x,y\right) \in T^{2}$.
\end{remark}

We say the point $(x,y)\in T^2$ is a {\bf regular point} of a function $f$, if all quadrant limits in (\ref{limits}) exist.

The well known Dirichlet-Jordan theorem (see \cite{Zy}) states that the
Fourier series of a function $g(x), \ x\in T$ of bounded variation converges
at every point $x$ to the value $\left[ g\left( x+0\right)
+g\left(x-0\right) \right] /2$. If $g$ is in addition continuous on $T$ the
Fourier series converges uniformly on $T$.

Hardy \cite{Ha} generalized the Dirichlet-Jordan theorem to the double
Fourier series. He proved that if function $f$ has bounded variation in
the sense of Hardy ($f\in BV$), then $S \left[ f\right] $ converges  to
$f^{\ast }(x,y)$ at any regular point $\left( x,y\right) $.
If $f$ is in addition continuous on $T^{2}$ then $S
\left[ f\right] $ converges uniformly on $T^{2}$.

\begin{Sa} [Sahakian \cite{Saha}] \label{Sa}
{The Fourier series of a function $f\in HBV$ converges to $f^\ast(x,y)$
in any regular point $\left( x,y\right) $. The convergence is uniform on any
compact $K\subset T^{2}$, where the function $f$ is continuous.}
\end{Sa}
Theorem S was proved in \cite{Saha} under assumption that the
function is continuous on some open set containing $K$, while O. Sargsyan
noticed in \cite{SG}, that the continuity of $f$ on the compact $K$ is
sufficient.
\begin{definition}
We say that the class of functions $V \subset L^{1}(T^{2})$ is a class of
convergence on $T^{2}$, if for any function $f\in V $

 1) the Fourier series of $f$ converges to $f^{\ast }({x,y})$ at any
regular point $(x,y)$,

2) the convergence is uniform on any compact $K\subset T^{2}$,  where the function $f$ is continuous.
\end{definition}

The following results immediately follow from Theorems \ref{T1}, \ref{T2}, Corollary \ref{cor1.2} and Theorem S.

\begin{theorem}[U. Goginava, A. Sahakian \cite{EJA}]\label{d=2}
Let $\Lambda =\{\lambda _{n}\}$ with $\lambda _{n}=n\gamma_n$ and $\gamma_{n}\geq
\gamma _{n+1}>0,\ n=1,2,....\,\,\,$. \newline
\indent
1) If
\begin{equation*}
\sum_{n=1}^{\infty }\frac{\gamma _{n}}{n}<\infty,
\end{equation*}
then the class  $P\Lambda BV$ is a class of convergence on $T^{2}$.\newline
\indent 2) If $\gamma_n=O(\gamma_{n^{[1+\delta]}})$ for some $\delta>0$ and
\begin{equation*}  \label{T1-11}
\sum_{n=1}^{\infty }\frac{\gamma _{n}}{n}=\infty,
\end{equation*}
then
then there exists a continuous function $f\in P\Lambda BV$, the Fourier
series of which diverges over cubes at $\left( 0,0\right) .$
\end{theorem}
\begin{theorem}[U. Goginava, A. Sahakian \cite{EJA}]
The set of functions $f$ satisfying
\begin{equation*}
\sum\limits_{n=1}^{\infty }\frac{\sqrt{v_{j}\left( n,f\right) }}{n^{3/2}}%
<\infty ,\quad j=1,2,
\end{equation*}
is a class of convergence on $T^{2}$.
\end{theorem}

\begin{corollary}
The set of functions $f$ satisfying $v_{1}\left(n,f\right) =O\left(n^{\alpha }\right)$, $ v_{2}\left(n,f\right) =O\left(n^{\beta }\right),$ $0<\alpha ,\beta <1$,
is a class of convergence on $T^{2}$.
\end{corollary}

\begin{theorem} [U. Goginava \cite{GoEJA}]
The class  $PBV_{p},\ p\geq 1$, is a class of convergence on $T^{2}$.
\end{theorem}

From Theorem \ref{d=2} follows that for any $\delta >0$ the class
 $f\in P\left\{ \frac{n}{\log ^{1+\delta }n}\right\} BV$ is a class of
convergence. Moreover, one can not take here $\delta =0$. It is interesting to
compare this result with the following one obtained by M. Dyachenko and D. Waterman in
\cite{DW}.

\begin{TDW}[M. Dyachenko and D. Waterman \cite{DW}]
\label{TDW} If $f\in \left\{\frac n {\log n}\right\}^*BV$, then in any point
$(x,y)\in T^2$ the quadrant limits (\ref{limits}) exist and the double Fourier
series of $f$ converges to $f^\ast(x,y)$.\newline
Moreover, the sequence $\left\{\frac n {\log n}\right\}$ can not be replaced
with any sequence $\left\{\frac {n\alpha_n} {\log n}\right\}$, where $%
\alpha_n\to \infty$.
\end{TDW}

It is easy to show \textrm{(see\cite{DW})}, that $\left\{\frac n {\log
n}\right\}^*BV\subset HBV$, hence the convergence part of Theorem DW follows
from Theorem S. It is essential that the condition $f\in \left\{\frac n
{\log n}\right\}^*BV$ guaranties the existence of quadrant limits.

The following theorem immediately follows from Theorem \ref{TT}
and Theorem S.

\begin{theorem}[U. Goginava, A. Sahakian \cite{GMJ}]\label{GS}
If $\Lambda =\{\lambda _{n}\}$  and
\begin{equation*}
{\limsup\limits_{n\rightarrow \infty }}\frac{\lambda _{n}\log n}{n}<\infty ,  \label{cond1}
\end{equation*}%
then the class $\Lambda^{\#}BV$ is a class of convergence on $T^{2}$.

In particular, the class $\left\{ \frac{n}{\log n}\right\} ^{\#}BV$ is a class of convergence on $T^{2}$.
\end{theorem}

Theorem DW and (\ref{cond0}) imply that the sequence $\left\{ \frac{n}{\log
n}\right\} $ in Theorem \ref{GS} can not be replaced with any sequence $%
\left\{ \frac{n\alpha _{n}}{\log n}\right\} $, where $\alpha _{n}\rightarrow
\infty $.

Theorems \ref{phi}, \ref{v(n)} and \ref{GS} imply

\begin{theorem}[U. Goginava, A. Sahakian \cite{GMJ}]
The class $B^{\#}V_{\Phi }$ is a class of convergence on $T^{2}$, provided that (\ref{T1-1}) and (\ref{Phi}) hold.
\end{theorem}

\begin{theorem}[U. Goginava, A. Sahakian \cite{GMJ}]
Let%
\begin{equation*}
\sum\limits_{n=1}^{\infty }\frac{v_{s}^{\#}\left( f,n\right) \log n }{n^{2}}<\infty ,~\ \ s=1,2.
\end{equation*}%
Then in any point $(x,y)\in T^{2}$ the quadrant limits (\ref{limits}) exist
and the double Fourier series of $\ f$ converges to $f^\ast(x,y)$. The convergence is uniform on any compact $K\in
T^{2}$, if $\ f$ is continuous on $K$.
\end{theorem}

\section{Ces\`{a}ro
Summability of double Fourier
series}
For one-dimensional Fourier series D. Waterman has proved the following theorem.
\begin{Wa2}[D. Waterman \cite{Wa2}]\label{DW2}
Let $0<\alpha <1$. The Fourier series of a function $f\in \{n^{1-\alpha
}\}BV $ is everywhere $\left( C,-\alpha \right) $ bounded and is uniformly $%
\left( C,-\alpha \right) $ bounded on each closed interval of continuity of $%
f$. \newline
\indent If $f\in C\{n^{1-\alpha }\}BV$, then $S[f]$ is everywhere $\left(
C,-\alpha \right) $ summable to the value $\left[ f\left( x+0\right)
+f\left( x-0\right) \right] /2$ and the summability is uniform on each
closed interval of continuity.
\end{Wa2}

Later A.  Sablin proved in \cite{Sab}, that for $0<\alpha<1$ the classes $%
\{n^{1-\alpha }\}BV$ and $C\{n^{1-\alpha }\}BV$ coincide.

For double Fourier series  the {\bf Ces\`{a}ro  $(C;\alpha ,\beta )$-means} 
 of a function $f\in L^1(T^2)$ are defined by
\begin{equation*}
\sigma _{n,m}^{\alpha ,\beta }(f;x,y):=\frac{1}{A_{n}^{\alpha }}\frac{1}{%
A_{m}^{\beta }}\sum_{i=0}^{n}\sum_{j=0}^{m}A_{n-i}^{\alpha -1}A_{m-j}^{\beta
-1}S_{i,j}\left[ f,(x,y)\right],
\end{equation*}
where $\alpha ,\beta >-1$ and 
\begin{equation*}
A_{0}^{\alpha }=1,\,\,\,A_{k}^{\alpha }=\frac{(\alpha +1)\cdots (\alpha +k)}{%
k!},\quad k=1,2,....
\end{equation*}
The double Fourier series of $f$ is said to be $\left( C;\alpha
,\beta \right) $  {\bf summable to $s$ in a point }$\left( x,y\right) $, if
$$
\lim_{n,m\to \infty}\sigma _{n,m}^{\alpha ,\beta }(f;x,y)=s.
$$

L. Zhizhiashvili  has investigated the convergence of Ces\`{a}ro
means of double Fourier series of functions of bounded variation. In particular, the following
theorem was proved.

\begin{Zh}[L. Zhizhiashvili \protect\cite{Zh}]
If $f\in BV$, then the double Fourier series of $f$ is $\left( C;-\alpha
,-\beta \right) $ summable to $f^\ast(x,y)$
in any regular point $\left( x,y\right) $. The convergence is uniform on any
compact $K$, where the function $f$ is continuous.
\end{Zh}

For functions of partial bounded variation the problem was considered by the
first author.
\begin{Go2}[{U. Goginava} \cite{GoPRI}] Let $\alpha>0,\ \beta >0$.

1) If  $\alpha +\beta <1$, then for any $f\in C\left( T^{2}\right) \cap PBV$
the double Fourier series of  $f$
is uniformly $(C;-\alpha ,-\beta )$ summable to $f$.

2) If $\alpha +\beta \geq 1$, then there exists a
continuous function $f_{0}\in PBV$ such that the sequence $\sigma _{n,n}^{-\alpha ,-\beta }\left( f_{0};0,0\right) \,$ diverges.
\end{Go2}

In \cite{Szeg} we consider the following problem. \textit{Let }$\alpha
,\beta \in \left( 0,1\right) ,\,\alpha +\beta <1.$ \textit{Under what
conditions on the sequence }$\Lambda =\{\lambda _{n}\}$ \textit{the double
Fourier series of any function $f\in P\Lambda BV$ is $(C;-\alpha ,-\beta )$
summable?}

\begin{theorem}[U. Goginava, A. Sahakian \cite{Szeg}]
Let $\alpha ,\beta \in \left( 0,1\right) ,\ \alpha +\beta <1$ and let the sequence $\Lambda =\{\lambda _{k}\}$ be such that ${\lambda _{k}}{k^{\left( \alpha +\beta \right)-1 }}\downarrow 0$. 

1) If
\begin{equation*}
\qquad
\sum\limits_{k=1}^{\infty }\frac{\lambda _{k}}{k^{2-\left( \alpha +\beta
\right) }}<\infty,
\end{equation*}%
then the double Fourier series of any function $f\in P\Lambda BV$ is $\left(
C;-\alpha ,-\beta \right) $ summable to $f^*(x,y)$ at any regular point $\left( x,y\right) $. The
summability is uniform on any compact $K$, if  $f$ is continuous on the neighborhood of $K$.

2) If 
\begin{equation*}
\sum\limits_{k=1}^{\infty }\frac{\lambda _{k}}{k^{2-\left( \alpha +\beta
\right) }}=\infty,
\end{equation*}%
then there exists a continuous function $f\in P\Lambda BV$ for which the  $\left(
C;-\alpha ,-\beta \right) $ means of the double Fourier series diverges
over cubes at $\left( 0,0\right) .$
\end{theorem}

\begin{corollary}[U. Goginava, A. Sahakian \cite{Szeg}]
Let $\alpha ,\beta \in \left( 0,1\right) ,\,\alpha +\beta <1$. \newline

1) If $f\in P\left\{ \frac{n^{1-\left( \alpha +\beta \right) }}{\log
^{1+\varepsilon } n }\right\} BV$ for some $\varepsilon >0$,
then the double Fourier series of the function $f$ is $\left( C;-\alpha
,-\beta \right) $ summable to  $f^*(x,y)$
in any regular point $\left( x,y\right) $. The summability is uniform on any compact $K$, if
 $f$ is continuous on the neighborhood of $K$.

2) There exists a continuous function $f\in P\left\{ \frac{n^{1-\left(
\alpha +\beta \right) }}{\log \left( n+1\right) }\right\} BV$ such that $%
\left( C;-\alpha ,-\beta \right) $ means of two-dimensional Fourier series
of $f$ diverges over cubes at $\left( 0,0\right) .$
\end{corollary}

\begin{corollary}[U. Goginava, A. Sahakian \cite{Szeg}]
Let $\alpha,\beta \in \left( 0,1\right) ,\,\alpha +\beta <1$ \newline
and $\,\,f\in PBV$. Then the double Fourier series of the
function $f$ is $\left( C;-\alpha ,-\beta \right) $ summable to $f^*(x,y)$
 in any regular point $\left( x,y\right) $. The
summability is uniform on any compact $K$, if
 $f$ is continuous on the neighborhood of $K$.
\end{corollary}

In \cite{SMJ} the following problem was considred. \textit{Let }$\alpha ,\beta
\in \left( 0,1\right) ,\,\alpha +\beta <1.$ \textit{Under what conditions on
the sequence }$\Lambda =\{\lambda _{n}\}$ \textit{the double Fourier series
of any function $f\in C\Lambda ^{\#}BV$ is $(C;-\alpha ,-\beta )$ summable.}

\begin{theorem}[U. Goginava \cite{SMJ}]
\label{Main}a) Let $\alpha ,\beta \in \left( 0,1\right) ,\,\alpha +\beta <1$
and $f\in C\left\{ n^{1-\left( \alpha +\beta \right) }\right\} ^{\#}BV$.
Then the double Fourier series of  $f$ is $\left( C;-\alpha ,-\beta \right) $ summable to $f^*(x,y)$
 in any point $\left( x,y\right) $. The
summability is uniform on any compact $K\subset \mathbb{T}^{2}$, if
 $f$ is continuous on the neighborhood of $K$.\newline
b) Let $\Lambda :=\left\{ n^{1-\left( \alpha +\beta \right) }\xi
_{n}\right\} $, where $\xi _{n}\uparrow \infty $ as $n\rightarrow \infty $.
Then \ there exists a function $f\in C\left( \mathbb{T}^{2}\right) \cap
C\Lambda ^{\#}V$ for which $\left( C;-\alpha ,-\beta \right) $-means of
double Fourier series diverges unboundedly at $\left( 0,0\right) $.
\end{theorem}


Theorems \ref{continous}, \ref{MV} and  \ref{Main} imply the following results.

\begin{theorem}
Let $\alpha ,\beta \in \left( 0,1\right) ,\,\alpha +\beta <1$ and $f\in
\left\{ n^{1-\left( \alpha +\beta \right) }\right\} ^{\#}BV$. Then the double Fourier series of  $f$ is $\left( C;-\alpha ,-\beta \right) $ summable to $f^*(x,y)$
 in any point $\left( x,y\right) $.  The
summability is uniform on any compact $K\subset \mathbb{T}^{2}$, if
 $f$ is continuous on the neighborhood of $K$.\newline
\end{theorem}


\begin{theorem}
Let $\alpha ,\beta \in \left( 0,1\right) ,\,\alpha +\beta <1$ and%
\begin{equation*}
\sum\limits_{j=1}^{\infty }\frac{v_{s}^{\#}\left( f;2^{j}\right) }{%
2^{j\left( 1-\left( \alpha +\beta \right) \right) }}<\infty ,\quad s=1,2.
\end{equation*}%
Then the double Fourier series of  $f$ is $\left( C;-\alpha ,-\beta \right) $ summable to $f^*(x,y)$
 in any point $\left( x,y\right) $. The
summability is uniform on any compact $K\subset \mathbb{T}^{2}$, if $f$ is continuous on the neighborhood of $K$.\newline
\end{theorem}

\section{Classes of Functions of $d$ variables of Bounded Generalized
Variation}

Consider a function $f\left( x\right) $ defined on the
d-dimensional cube $T^{d}$ and a collection
of intervals
\begin{equation*}
J^{k}=\left( a^{k},b^{k}\right) \subset T,\qquad k=1,2,\ldots d.
\end{equation*}
For $d=1$ we set
\begin{equation*}
f\left( J^{1}\right) :=f\left( b^{1}\right) -f\left( a^{1}\right) .
\end{equation*}%
If for any function of $d-1$ variables the expression $f\left( J^{1}\times
\cdots \times J^{d-1}\right) $ is already defined, then for a function $f$ of $d$
variables the {\bf mixed difference} is defined as follows:
\begin{equation*}
f\left( J^{1}\times \cdots \times J^{d}\right) :=f\left( J^{1}\times \cdots
\times J^{d-1},b^{d}\right) -f\left( J^{1}\times \cdots \times
J^{d-1},a^{d}\right) .
\end{equation*}
For sequences of positive numbers
\begin{equation*}
\Lambda ^{j}=\{\lambda _{n}^{j}\}_{n=1}^{\infty },\quad
\lim_{n\to\infty}\lambda^j_n=\infty,\quad j=1,2,\ldots ,d,
\end{equation*}
and for a function $f(x)$, $x=(x_1,\ldots,x_d)\in T^d$ the $\left( \Lambda
^{1},\ldots ,\Lambda ^{d}\right) $-{\bf variation of $f$ with respect to
the index set }$D:=\{1,2,...,d\}$ is defined as follows:
\begin{equation*}
\left\{ \Lambda ^{1},\ldots ,\Lambda ^{d}\right\} V^{D}\left( f,T^{d}\right)
:=\sup\limits_{\{I_{i_{j}}^{j}\}\in \Omega }\ \sum\limits_{i_{1},...,i_{d}}%
\frac{\left\vert f\left( I_{i_{1}}^{1}\times \cdots \times
I_{i_{d}}^{d}\right) \right\vert }{\lambda^1 _{i_{1}}\cdots \lambda^d
_{i_{d}}}.
\end{equation*}

For an index set $\alpha =\{j_{1},...,j_{p}\}\subset D$ and any $x=\left(
x_{1},...,x_{d}\right) \in R^{d}$ we set ${\widetilde{\alpha }}:=D\setminus
\alpha $ and denote by $x_{\alpha }$ the vector of $R^{p}$ consisting of
components $x_{j},j\in \alpha $, i.e.
\begin{equation*}
x_{\alpha }=\left( x_{j_{1}},...,x_{j_{p}}\right) \in R^{p}.
\end{equation*}

By
\begin{equation*}
\left\{ \Lambda ^{j_{1}},...,\Lambda ^{j_{p}}\right\} V^{{\alpha }}\left(
f,x_{\widetilde{\alpha }},T^d\right) \quad \text {and}\quad f\left(
I_{i_{j_{1}}}^{1}\times \cdots \times I_{i_{j_{p}}}^{p},x_{\widetilde{\alpha
}}\right)
\end{equation*}
we denote respectively the $\left( \Lambda ^{j_{1}},...,\Lambda
^{j_{p}}\right) $-variation over the $p$-dimensional cube $T^{p}$ and mixed
difference of $f$ as a function of variables $x_{^{j_{1}}},...,x_{j_{p}}$
with fixed values $x_{^{\widetilde{\alpha }}}$ of other variables. The $%
\left( \Lambda ^{j_{1}},...,\Lambda ^{j_{p}}\right) ${\bf-variation of $%
f$\ with respect to the index set} ${\alpha }$ is defined as
follows:
\begin{equation*}
\left\{ \Lambda ^{j_{1}},...,\Lambda ^{j_{p}}\right\} V^{{\alpha }}\left(
f,T^{p}\right) =\sup\limits_{x_{^{{\widetilde{\alpha }}}}\in T^{d-p}}\left\{{%
\Lambda ^{j_{1}},...,\Lambda ^{j_{p}}}\right\}V^{{\alpha }}\left( f,x_{^{%
\widetilde{\alpha }}},T^{d}\right) .
\end{equation*}

\begin{definition}
We say that the function $f$ has total Bounded $\left( \Lambda
^{1},...,\Lambda ^{d}\right) $-variation on $T^{d}$ and write $f\in \left\{
\Lambda ^{1},...,\Lambda ^{d}\right\}BV\left( T^{d}\right) $, if
\begin{equation*}
\left\{ \Lambda ^{1},...,\Lambda ^{d}\right\}
V(f,T^{d}):=\sum\limits_{\alpha \subset D}\left\{ \Lambda ^{1},...,\Lambda
^{d}\right\}V^{{\alpha }}\left( f,T^{d}\right) <\infty .
\end{equation*}
\end{definition}

\begin{definition}
We say that the function $f$ is continuous in $\left( \Lambda
^{1},...,\Lambda ^{d}\right) $-variation on $T^{d}$ and write $f\in C\left\{
\Lambda ^{1},...,\Lambda ^{d}\right\} V\left( T^{d}\right) $, if%
\begin{equation*}
\lim\limits_{n\rightarrow \infty }\left\{ \Lambda ^{j_{1}},...,\Lambda
^{j_{k-1}},\Lambda _{n}^{j_{k}},\Lambda ^{j_{k+1}},...,\Lambda
^{j_{p}}\right\} V^{{\alpha }}\left( f,T^{d}\right) =0,\qquad k=1,2,\ldots ,p
\end{equation*}%
for any $\alpha \subset D,\ \alpha :=\{j_{1},...,j_{p}\}$, where $\Lambda
_{n}^{j_{k}}:=\left\{ \lambda _{s}^{j_{k}}\right\} _{s=n}^{\infty }$.
\end{definition}
The continuity of a function in $\Lambda$-variation was introduced by D. Waterman \cite{Wa2}
and was investigated in details by A. Bakhvalov (see \cite{Bakh}, \cite{Bakh2} and references therein).
This property is important for applications in the theory of Fourier series (see Theorem B1 in Section 5).
\begin{definition}
We say that the function $f$ has Bounded Partial $\left( \Lambda
^{1},...,\Lambda ^{d}\right) $-variation and write $f\in P\left\{ \Lambda
^{1},...,\Lambda ^{d}\right\} BV\left( T^{d}\right) $ if
\begin{equation*}
P\left\{ \Lambda ^{1},...,\Lambda ^{d}\right\}
V(f,T^{d}):=\sum\limits_{i=1}^{d}\Lambda ^{i}V^{\{i\}}\left( f,T^{d}\right)
<\infty .
\end{equation*}
\end{definition}
In the case when $\Lambda ^{1}=\cdots =\Lambda ^{d}=\Lambda $ we set
\begin{eqnarray*}
\Lambda BV(T^{d}):= &&\{\Lambda ^{1},...,\Lambda ^{d}\}BV(T^{d}), \\
C\Lambda V(T^{d}):= &&C\{\Lambda ^{1},...,\Lambda ^{d}\}V(T^{d}), \\
P\Lambda BV(T^{d}):= &&P\{\Lambda ^{1},...,\Lambda ^{d}\}BV(T^{d}).
\end{eqnarray*}%
If $\lambda _{n}=n$ for all $n=1,2\ldots $ we say \textit{Harmonic
Variation} instead of $\Lambda $-variation and write $H$ instead of $%
\Lambda, i.e. $ $HBV$, $PHBV$, $CHV$, ets.

\begin{theorem}[U. Goginava, A. Sahakian \cite{AM}] \label{41} 
Let $\Lambda =\{\lambda _{n}\}_{n=1}^{\infty }$ and $d\geq 2$. If
${\lambda _{n}}/{n}\downarrow 0$ and
\begin{equation*}\label{lambda}
\sum\limits_{n=1}^{\infty }\frac{\lambda _{n}\log ^{d-2}n}{n^{2}}<\infty ,
\end{equation*}%
then $P\Lambda BV(T^{d})\subset CHV(T^{d})$.
\end{theorem}
For a sequence $\Lambda =\{\lambda _{n}\}_{n=1}^{\infty }$ we denote%
\begin{equation*}
\Lambda ^{\#}V_{s}\left( f,T^{d}\right) :=\sup\limits_{\left\{ x^{i}{\left\{
s\right\} }\right\} \subset T^{d-1}}\sup\limits_{\left\{ I_{i}^{s}\right\}
\in \Omega }\sum\limits_{i}\frac{\left\vert f\left( I_{i}^{s},x^{i}{\left\{
s\right\} }\right) \right\vert }{\lambda _{i}},
\end{equation*}%
where
\begin{equation*}
x^{i}{\left\{ s\right\} }:=\left( x_{1}^{i},\ldots
,x_{s-1}^{i},x_{s+1}^{i},\ldots ,x_{d}^{i}\right) \quad \text{for}\quad
x^{i}:=\left( x_{1}^{i},\ldots ,x_{d}^{i}\right) .  \label{xis}
\end{equation*}

\begin{definition}
\label{def5} We say that  $f\in \Lambda
^{\#}BV\left( T^{d}\right) $, if%
\begin{equation*}
\Lambda ^{\#}V\left( f,T^{d}\right) :=\sum\limits_{s=1}^{d}\Lambda
^{\#}V_{s}\left( f,T^{d}\right) <\infty .
\end{equation*}
\end{definition}


\begin{theorem}[U. Goginava, A. Sahakian \cite{SBM}]
If $\Lambda =\left\{ \lambda _{n}\right\}$ with
\begin{equation*}\label{1}
\lambda _{n}=%
\frac{n}{\log ^{d-1}n },\quad n=2,3,\ldots,
\end{equation*}%
then $\Lambda
^{\#}BV(T^{d})\subset HBV(T^{d})$.
\end{theorem}

Now, we denote
\begin{equation*}
\Delta :=\{\delta =(\delta _{1},\ldots ,\delta _{d}):\delta _{i}=\pm 1,\
i=1,2,\ldots ,d\}  \label{Delta}
\end{equation*}%
and
\begin{equation*}
\pi _{\varepsilon \delta }(x):=(x_{1},\,x_{1}+\varepsilon \delta _{1})\times
\cdots \times (x_{d},\,x_{d}+\varepsilon \delta _{d}),
\end{equation*}%
for $x=(x_{1},\ldots ,x_{d})\in R^{d}$ and $\varepsilon >0$. We set $\pi
_{\delta }(x):=\pi _{\varepsilon \delta }(x)$, if $\varepsilon =1$.

For a function $f$ and $\delta\in \Delta$ we set
\begin{equation}  \label{lim}
f_\delta(x):=\lim_{t\in \pi_\delta(x),\ t\to x} f(t),
\end{equation}
if the last limit exists.

\begin{theorem}[U. Goginava, A. Sahakian \cite{SBM}]
\label{th2} Suppose  $\Lambda=\{\lambda_n\}$ and $f\in \Lambda ^{\#}BV\left( T^{d}\right)$.

a) If the limit $f_\delta(x)$ exists for some $x=(x_1,\ldots,x_d)\in T^d$
and some $\delta=(\delta_1,\ldots,\delta_d)\in \Delta$, then
\begin{equation*}  \label{th2a}
\lim\limits_{\varepsilon \rightarrow 0}\Lambda ^{\#}V\left(
f,\pi_{\varepsilon\delta}(x)\right) =0.
\end{equation*}

b) If $f$ is continuous on some compact $K\subset T^{d}$, then
\begin{equation*}
\lim_{\varepsilon \rightarrow 0}\Lambda ^{\#}V\left( f,\left[
x_{1}-\varepsilon ,x_{1}+\varepsilon \right] \times \cdots \times \left[
x_{d}-\varepsilon ,x_{d}+\varepsilon \right] \right) =0  \label{th2b}
\end{equation*}%
uniformly with respect to $x=(x_{1},\ldots ,x_{d})\in K$.
\end{theorem}

\begin{theorem}[U. Goginava, A. Sahakian \cite{SBM}]
If the function $f(x),\ x\in T^{d}$ satisfies the condition
\begin{equation*}
\sum\limits_{n=1}^{\infty }\frac{v_{s}^{\#}\left( f,n\right) \log
^{d-1}n }{n^{2}}<\infty ,\qquad \ s=1,2,...,d,
\end{equation*}%
then $f\in \left\{ \frac{n}{\log ^{d-1}n }\right\}
^{\#}BV\left( T^{d}\right) .$
\end{theorem}

\section{\protect\medskip Convergence of multiple Fourier series}

The Fourier series of function $f\in L^{1}\left( T^{d}\right) $ with respect
to the trigonometric system is the series
\begin{equation*}
S\left[ f\right] :=\sum_{n_{1},...,n_{d}=-\infty }^{+\infty }\widehat{f}%
\left( n_{1},....,n_{d}\right) e^{i\left( n_{1}x+\cdots +n_{d}x_{d}\right) },
\end{equation*}%
where
\begin{equation*}
\widehat{f}\left( n_{1},....,n_{d}\right) =\frac{1}{\left( 2\pi \right) ^{d}}%
\int_{T^{d}}f(x_{1},...,x_{d})e^{-i\left( n_{1}x_{1}+\cdots
+n_{d}x_{d}\right) }dx_{1}\cdots dx_{d}
\end{equation*}%
are the Fourier coefficients of $f$. The rectangular partial sums are
defined as follows:
$$
S_{N_{1},...,N_{d}}\left [f, (x_{1},...,x_{d})\right] =\sum_{n_{1}=-N_{1}}^{N_{1}}\cdots \sum_{n_{d}=-N_{d}}^{N_{d}}\widehat{f}%
\left( n_{1},....,n_{d}\right) e^{i\left( n_{1}x_{1}+\cdots
+n_{d}x_{d}\right) }
$$
We denote by $C(T^{d})$ the space of continuous and $2\pi $-periodic with
respect to each variable functions with the norm
\begin{equation*}
\Vert f\Vert _{C}:=\sup_{\left( x^{1},\ldots,\,x^{d}\right) \in
T^{d}}|f(x^{1},\ldots,x^{d})|.
\end{equation*}

We say that the point $x:=\left( x^{1},\ldots ,x^{d}\right) \in T^d$ is a {\bf
regular point} of a function $f$ if the limits (\ref{lim}) exist for all $\delta\in \Delta$.
For a regular point $x\in T^d$ we denote%
\begin{equation*}
f^{\ast }\left( x\right) :=\frac{1}{2^{d}}\sum_{\delta\in \Delta} f_\delta(x).  \label{limit}
\end{equation*}

\begin{definition}
We say that the class of functions $V \subset L^{1}(T^{d})$ is a class of
convergence on $T^{d}$, if for any function $f\in V $

1) the Fourier series of $f$ converges to $f^{\ast }({x})$ at any regular
point ${x}\in T^{d}$,

2) the convergence is uniform on any compact $K\subset T^{d}$, if $f$ is
continuous on the neighborhood of $K$.
\end{definition}
In \cite{Bakh} A. Bakhvalov showed that the class $HBV(T^{d})$ is not a class of
convergence on $T^{d}$, if $d>2$. On the other hand, he proved the following

\begin{Bakh1}[A. Bakhvalov \protect\cite{Bakh}]
\label{B} The class $CHV(T^{d})$ is a class of convergence on $T^{d}$ for any $d=1,2,\ldots $
\end{Bakh1}

Convergence of spherical and other partial sums of d-dimensional Fourier
series of functions of bounded $\Lambda $-variation was investigated in
deatails by M. Dyachenko \cite{D1,D2}, A. Bakhvalov \cite{Bakh,Bakh3}.

The first part of the next theorem is a consequence of Theorem \ref{41} and Theorem B1.

\begin{theorem}[[U. Goginava, A. Sahakian \cite{AM}]\label{t30}
Let $\Lambda =\{\lambda _{n}\}$ and $d\geq 2$.

a) If $\lambda_n/n \downarrow 0$ and
\begin{equation*}
\sum\limits_{n=1}^{\infty }\frac{\lambda _{n}\log ^{d-2}n}{n^{2}}<\infty ,
\end{equation*}%
then $P\Lambda BV$ is a class of convergence on $T^{d}$.

\medskip b) If
$ \frac{\lambda _{n}}{n}=O\left( \frac{\lambda _{\lbrack n^{\delta }]}}{%
[n^{\delta }]}\right)$
for some $\delta >1$, and
\begin{equation*}
\sum\limits_{n=1}^{\infty }\frac{\lambda _{n}\log ^{d-2}n}{n^{2}}=\infty ,
\end{equation*}%
then there exists a continuous function $f\in P\Lambda BV$, the Fourier
series of which diverges at $\left( 0,\ldots ,0\right) .$
\end{theorem}

Theorem \ref{t30} imply

\begin{corollary}
\label{c3} a) If $\Lambda =\left\{ \lambda _{n}\right\} _{n=1}^{\infty }$
with
\begin{equation*}
\lambda _{n}=\frac{n}{\log ^{d-1+\varepsilon }n},\qquad n=2,3,\ldots
\end{equation*}%
for some $\varepsilon >0$, then the class $P\Lambda BV$ is a class of
convergence on $T^{d}$.

\medskip b) If $\Lambda =\left\{ \lambda _{n}\right\} _{n=1}^{\infty }$ with
\begin{equation*}
\lambda _{n}=\frac{n}{\log ^{d-1}n},\qquad n=2,3,\ldots ,
\end{equation*}%
then the class $P\Lambda BV$ is not a class of convergence on $T^{d}$.
\end{corollary}

\begin{theorem}[Goginava, Sahakian \cite{SBM}]
\label{main} a) If $\Lambda =\left\{ \lambda _{n}\right\} _{n=1}^{\infty }$
with
\begin{equation*}
\lambda _{n}=\frac{n}{\log ^{d-1}n},\qquad n=2,3,\ldots,
\end{equation*}%
then the class $\Lambda ^{\#}BV\left( T^{d}\right) $ is a class of
convergence on $T^{d}$.

b)If $\Lambda =\left\{ \lambda _{n}\right\} _{n=1}^{\infty }$ with
\begin{equation*}
\lambda _{n}:=\left\{ \frac{n\xi _{n}}{\log ^{d-1} n}%
\right\} ,\quad n=2,3,\ldots,
\end{equation*}%
where $\xi _{n}\rightarrow \infty $ as $n\rightarrow \infty $, then there
exists a continuous function $f\in \Lambda ^{\#}BV\left( T^{d}\right) $ such
that the cubical partial sums of $d$-dimensional Fourier series of $f$
diverge unboundedly at $\left( 0,...,0\right) \in T^{d}$.
\end{theorem}

\begin{theorem}[Goginava, Sahakian \cite{SBM}]
For any $d>1$ the class of functions $f(x),\ x\in T^{d}$ satisfying the
following condition
\begin{equation*}
\sum\limits_{n=1}^{\infty }\frac{v_{s}^{\#}\left( f,n\right) \log
^{d-1}n }{n^{2}}<\infty ,~\ \ s=1,...,d,
\end{equation*}%
is a class of convergence.
\end{theorem}

\section{\protect Ces\`{a}ro summability of $d$-dimensional Fourier series}

The Ces\`{a}ro $\left( C;{\alpha }_{1},...,\alpha _{d}\right) \,$means of $d$%
-dimensional Fourier series of function $f\in L^1(T^d)$ is defined by
\begin{multline*}
\sigma _{m_{1},...,m_{d}}^{{\alpha }_{1},...,\alpha
_{d}}[f;(x_{1},...,x_{d})] \\
:=\left( \prod\limits_{i=1}^{d}A_{m_{i}}^{\alpha _{i}}\right)
^{-1}\sum\limits_{p_{1}=0}^{m_{1}}\cdots
\sum\limits_{p_{d}=0}^{m_{d}}\prod\limits_{i=1}^{d}A_{m_{i}-p_{i}}^{\alpha
_{i}-1}S_{p_{1},...,p_{d}}[f,(x_{1},...,x_{d})\mathbf{]},
\end{multline*}%
where{\large \ }
\begin{equation*}
A_{0}^{\alpha }=1,\,\,A_{n}^{\alpha }=\frac{\left( \alpha +1\right) \cdots
\left( \alpha +n\right) }{n!},\qquad \alpha>-1.
\end{equation*}
The Fourier series $S[f]$ is said to be
$\left( C;-\alpha _{1},...,-\alpha _{d}\right) $ summable to $s$
in a point $\left( x_1,\ldots,x_d\right)$, if
$$
\sigma _{m_{1},...,m_{d}}^{{\alpha }_{1},...,\alpha
_{d}}[f;(x_{1},...,x_{d})]\to s\quad\text{as}\quad x_{1},...,x_{d}\to \infty.
$$
\begin{definition}
We say that the class of functions $\Omega \subset L^{1}(T^{d})$ is a class
of $\left( C;-\alpha _{1},...,-\alpha _{d}\right) $ summability on $T^{d}$,
if the Cesaro $\left( C;{-\alpha }_{1},...,-\alpha _{d}\right) \,$ means of
Fourier series of any function $f\in \Omega $ converges to $f^{\ast }({x})$
at any regular point ${x}\in T^{d}$. The summability is uniform on any
compact $K\subset T^{d}$, if in addition, $f$ is continuous on the
neighborhood of $K$.
\end{definition}
The multivariate analog of Theorem W2 from Section 3 was proved by A. Bakhvalov in \cite{Bakh2}.
\begin{Bakh2}[A. Bakhvalov  \cite{Bakh2}] For any numbers $\alpha _{1},...,\alpha _{d}\in \left( 0,1\right)$ the class $C\{n^{1-\alpha_1}\},\ldots \{n^{1-\alpha_d}\}V(T^d)$ is a class of $\left( C;-\alpha _{1},...,-\alpha _{d}\right) $
summabi\-lity on $T^{d}$.
\end{Bakh2}

In the next theorem we consider the problem of  $\left( C;-\alpha _{1},...,-\alpha _{d}\right) $ summability of the Fourier series of functions of bounded partial $\Lambda$-variation.
\begin{theorem}[U. Goginava, A. Sahakian \cite{Stek}]\label{t3}
Suppose $\alpha _{1},...,\alpha _{d}\in \left( 0,1\right)$, $\alpha _{1}+\cdots
+\alpha _{d}<1$ and the sequence $\Lambda =\{\lambda _{n}\}_{n=1}^{\infty }$ is such that
\begin{equation*}
\frac{\lambda _{n}}{n^{1-\left( \alpha _{1}+\cdots +\alpha _{d}\right) }}%
\downarrow 0\text{ .}
\end{equation*}

a) If
\begin{equation*}
\sum\limits_{n=1}^{\infty }\frac{\lambda _{n}}{n^{2-\left( \alpha
_{1}+\cdots +\alpha _{d}\right) }}<\infty ,  \label{lambda1}
\end{equation*}%
then $P\Lambda BV(T^d)$ is a class of $\left( C;-\alpha _{1},...,-\alpha
_{d}\right) $ summability on $T^{d}$.

\medskip b) If
\begin{equation*}
\sum\limits_{n=1}^{\infty }\frac{\lambda _{n}}{n^{2-\left( \alpha
_{1}+\cdots +\alpha _{d}\right) }}=\infty ,  \label{lambda3}
\end{equation*}%
then there exists a continuous function $f\in P\Lambda BV(T^d)$ for which the sequence  $\sigma
_{N,...,N}^{-\alpha _{1},...,-\alpha _{d}}[f,\left( 0,...,0\right) ]$
diverges.
\end{theorem}

\begin{corollary}[U. Goginava, A. Sahakian \cite{Stek}]
Suppose $\alpha _{1},...,\alpha _{d}\in \left( 0,1\right),$ $\alpha _{1}+\cdots
+\alpha _{d}<1$ and $\Lambda =\left\{ \lambda _{n}\right\} _{n=1}^{\infty }$.

a) If
\begin{equation*}
\lambda _{n}=\frac{n^{1-\left( \alpha _{1}+\cdots +\alpha _{d}\right) }}{%
\log ^{1+\varepsilon }n },\qquad n=2.3.\ldots
\end{equation*}%
for some $\varepsilon >0$, then the class $P\Lambda BV(T^d)$ is a class of $%
\left( C;-\alpha _{1},...,-\alpha _{d}\right) $ summability on $T^{d}$.

b) If
\begin{equation*}
\lambda _{n}=\frac{n^{1-\left( \alpha _{1}+\cdots +\alpha _{d}\right) }}{%
\log n },\qquad n=2.3.\ldots,
\end{equation*}%
then $P\Lambda BV(T^d)$ is not a class of $\left( C;-\alpha
_{1},...,-\alpha _{d}\right) $ summability on $T^{d}$.
\end{corollary}

\begin{theorem}[U. Goginava, A. Sahakian  \cite{Stek}]
Let $\alpha _{1},...,\alpha _{d}\in \left( 0,1\right),$ $\alpha _{1}+\cdots
+\alpha _{d}<1.$ Then the set of functions $f$ satisfying the conditions
\begin{equation*}
\sum\limits_{j=0}^{\infty }\frac{\left( v_{i}\left(
2^{j},f\right) \right) ^{\alpha _{i}/\left( \alpha _{1}+\cdots +\alpha
_{d}\right) }}{2^{j\left( \alpha _{i}/\left( \alpha _{1}+\cdots +\alpha
_{d}\right) -\alpha _{i}\right) }}<\infty\quad \text{for}\quad i=1,...,d,
\end{equation*}%
is a class of $\left( C;-\alpha _{1},...,-\alpha _{d}\right) $ summability
on $T^{d}$.
\end{theorem}

\begin{theorem}[U. Goginava, A. Sahakian  \cite{Stek}]
Suppose  $\alpha _{1},...,\alpha _{d}\in \left( 0,1\right)$, $\alpha _{1}+\cdots
+\alpha _{d}<1/p$, $\, p\geq 1.$ Then the class $PBV_{p}$ is a class of $\left(
C;-\alpha _{1},...,-\alpha _{d}\right) $ summability on $T^{d}$.
\end{theorem}

In \cite{GoPRI} the first author has proved that the class $PBV_{p}$ is not a
class of $\left( C;-\alpha _{1},...,-\alpha _{d}\right) $ summability on $%
T^{d}$, if  $\alpha _{1},...,\alpha _{d}\in \left( 0,1\right),$ and $\alpha
_{1}+\cdots +\alpha _{d}\geq 1/p.$

\begin{corollary}[U. Goginava, A. Sahakian  \cite{Stek}]
Suppose  $\alpha _{1},...,\alpha _{d}\in \left( 0,1\right)$, $\alpha _{1}+\cdots
+\alpha _{d}<1$. Then the set of functions $f$ satisfying 
\begin{equation*}
v_{i}\left( 2^{j},f\right) =O\left( 2^{j\gamma }\right)\quad \text{for}\quad i=1,...,d,
,\qquad
\end{equation*}%
 is a class of $\left( C;-\alpha _{1},...,-\alpha _{d}\right) $ summability
on $T^{d}$.
\end{corollary}

\end{document}